A rancher is planning to put up 220 yards of fencing. In the morning she puts up 80 yards, and in the afternoon she puts up 40% of the remaining fence. What percent of the fence did she put up that day.
Answer by nyc_function(2741)

220 - 80 = 140
140 * 0.40 = 56
On that day, she put up 140 + 56 yards or 196 yards of fence.
There are 220 yards in total.
Let p = percent of the fence that she put up on that day
p = 196/220
p = 89%